OpenStudy (anonymous):

2x-3y=9 Find the x- and y-intercepts and graph the equation plotting the intercepts. Please show all of your work.

OpenStudy (mertsj):

Find the x intercept by letting y = 0 Find the y intercept by letting x = 0

OpenStudy (anonymous):

To find the intercepts, put the equation into slope-intercept form: y=mx+b Subtract 2x from both sides:\[-3y=9-2x\]Divide both sides by -3:\[y = \frac{2}{3}x-3\] So now you can see that the slope is 2/3 and the y-intercept is -3 (which is what you would get for y if you let x=0. To find the x intercept, solve for x and then let y=0:\[x = \frac{3}{2}(y+3)\]So if y=0, x=9/2